OpenStudy (anonymous):

Apply the chain rule, in Leibniz notation, to determine dy/dx at the indicated value of x. a) y=u^2 +3u, u=Square root of x-1, x=5.

OpenStudy (amistre64):

you can determine dy/du ; and du/dx so multiply them together to get dy/dx

OpenStudy (anonymous):

dy/dx = dy/du * du/dx = 2u+3 * 1/2(1/sqrt(x-1)) = 71/4 or -1/4

OpenStudy (anonymous):

Dhatraditya when you do that, you get 7 as your answer...

OpenStudy (anonymous):

you should get 7/4 or 1/4 as the answer

OpenStudy (anonymous):

you would get this \[(2(\sqrt{5-1)}+3(1/2(5-1)^{1/2}\]

OpenStudy (anonymous):

sorry it is \[\frac{2\sqrt{x-1}+3}{2\sqrt{x-1}}\]

OpenStudy (anonymous):

sqrt 4 gives two values, +2 and -2 with +2 you get 7/4 with -2 you get -1/-4 = 1/4

OpenStudy (anonymous):

Ahhh... okay, Thanks. such a fail on my part :P

OpenStudy (anonymous):

you are welcome